# Physics-Informed Teleconnection-Aware Transformer for Global Subseasonal-to-Seasonal Forecasting

### v2

@@ 3.2 Multi-Scale Physics-Informed Neural ODE
To incorporate physical prior knowledge of atmospheric dynamics kashinath2021physics , we evolve each frequency band using a physics-informed neural ODE. Drawing from fluid dynamics principles, we formulate an ODE in the latent space vermaclimode that combines fundamental atmospheric processes with data-driven corrections to handle complex, unresolved dynamics.
@@
For each multi-scale component $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{X}^{(\ell)}=[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{H}]\in\mathbb{R}^{H\times D_{emb}}$, we define an ODE that evolves the latent state through time:
@@
$\frac{d\mathbf{x}_{i}}{dt}=\gamma\cdot\tanh\Bigg{(}{\color[rgb]{0.0078125,0.46484375,0.7421875}\underbrace{\boldsymbol{\nu}\odot(\mathbf{x}_{i+1}-2\mathbf{x}_{i}+\mathbf{x}_{i-1})}_{\smash{\text{diffusion}}}}+{\color[rgb]{0.84765625,0.26171875,0.08203125}\underbrace{\boldsymbol{\mu}\odot\frac{\mathbf{x}_{i+1}-\mathbf{x}_{i-1}}{2}}_{\smash{\text{advection}}}}+{\color[rgb]{0.1875,0.24609375,0.625}\underbrace{\boldsymbol{f}}_{\smash{\text{forcing}}}}+{\color[rgb]{0,0.5390625,0.48046875}\underbrace{\alpha\cdot\text{MLP}(\mathbf{x}_{i})}_{\smash{\text{neural correction}}}}\Bigg{)}$ (6)
@@
  • Diffusion term: $\color[rgb]{0.171875,0.2421875,0.3125}{\color[rgb]{0.0078125,0.46484375,0.7421875}\boldsymbol{\nu}\odot(\mathbf{x}_{i+1}-2\mathbf{x}_{i}+\mathbf{x}_{i-1})}$ approximates the second-order diffusion operator along latitude, representing the meridional diffusion of atmospheric properties. The learnable parameter $\color[rgb]{0.171875,0.2421875,0.3125}\boldsymbol{\nu}\in\mathbb{R}^{D_{emb}}$ modulates diffusion strength for each feature dimension.
@@
  • Advection term: $\color[rgb]{0.171875,0.2421875,0.3125}{\color[rgb]{0.84765625,0.26171875,0.08203125}\boldsymbol{\mu}\odot\frac{\mathbf{x}_{i+1}-\mathbf{x}_{i-1}}{2}}$ approximates the first-order spatial derivative, representing meridional transport by mean flow. The parameter $\color[rgb]{0.171875,0.2421875,0.3125}\boldsymbol{\mu}\in\mathbb{R}^{D_{emb}}$ controls advection intensity.
@@
  • Forcing term: $\color[rgb]{0.171875,0.2421875,0.3125}{\color[rgb]{0.1875,0.24609375,0.625}\boldsymbol{f}\in\mathbb{R}^{D_{emb}}}$ represents external forcing, analogous to radiative and diabatic processes in the climate system.
@@
  • Neural correction: $\color[rgb]{0.171875,0.2421875,0.3125}{\color[rgb]{0,0.5390625,0.48046875}\alpha\cdot\text{MLP}(\mathbf{x}_{i})}$ learns complex, nonlinear dynamics not captured by simplified physical terms. Parameter $\color[rgb]{0.171875,0.2421875,0.3125}\alpha$ balances the physical terms with data-driven corrections.
@@
For numerical stability and physical realism, we apply boundary conditions $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{x}_{0}=\mathbf{x}_{H+1}=0$ (zero-padding at the poles) and constrain the rate of change using a hyperbolic tangent function. $\color[rgb]{0.171875,0.2421875,0.3125}\gamma$ is a small scaling factor that moderates the rate of change to prevent numerical instabilities. We employ the Euler method with a fixed step size $\color[rgb]{0.171875,0.2421875,0.3125}\Delta t$: $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{x}_{i}(t+\Delta t)=\mathbf{x}_{i}(t)+\Delta t\cdot\frac{d\mathbf{x}_{i}}{dt}\Big{|}_{t}$. We evolve each frequency band $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{X}^{(\ell)}$ independently from time $\color[rgb]{0.171875,0.2421875,0.3125}t=0$ to $\color[rgb]{0.171875,0.2421875,0.3125}t=T$ (where $\color[rgb]{0.171875,0.2421875,0.3125}T$ represents one forecast step in latent space), yielding evolved representations $\color[rgb]{0.171875,0.2421875,0.3125}\tilde{\mathbf{X}}^{(\ell)}$. This scale-dependent evolution reflects physical reality that atmospheric processes operate differently across scales, diffusion dominates at smaller scales while advection plays a more significant role at larger scales vallis2017atmospheric .
3.3 Teleconnection-Aware Transformer
Teleconnection-Aware Self-Attention
@@ 3.3 Teleconnection-Aware Transformer
Teleconnections, statistically significant correlations between weather and climate events at distant locations, such as the North Atlantic Oscillation (NAO) and Madden-Julian Oscillation (MJO), are fundamental for accurate S2S forecasting wallace1981teleconnections . We encode these global patterns through a novel teleconnection-aware self-attention mechanism.
@@
For each frequency band $\color[rgb]{0.171875,0.2421875,0.3125}\tilde{\mathbf{X}}^{(\ell)}\in\mathbb{R}^{H\times D_{emb}}$, we first compute standard queries, keys, and values: $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{Q}=\tilde{\mathbf{X}}^{(\ell)}\mathbf{W}^{Q}$, $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{K}=\tilde{\mathbf{X}}^{(\ell)}\mathbf{W}^{K}$, $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{V}=\tilde{\mathbf{X}}^{(\ell)}\mathbf{W}^{V}$, where $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{W}^{Q},\mathbf{W}^{K},\mathbf{W}^{V}\in\mathbb{R}^{D_{emb}\times d_{k}}$ are learnable projections and $\color[rgb]{0.171875,0.2421875,0.3125}d_{k}=D_{emb}/N_{h}$ for $\color[rgb]{0.171875,0.2421875,0.3125}N_{h}$ heads.
@@
$\mathbf{c}=\sum_{j=1}^{n_{p}}\omega_{j}\mathbf{P}_{j},\quad\text{with}\quad\boldsymbol{\omega}=\text{softmax}(\mathbf{\bar{x}}\mathbf{W}^{p})\in\mathbb{R}^{n_{p}},$ (7)
where $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{\bar{x}}=\frac{1}{H}\sum_{i=1}^{H}\tilde{\mathbf{x}}_{i}$ is the global atmosphere state, $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{W}^{p}\in\mathbb{R}^{D_{emb}\times n_{p}}$ projects this state onto $\color[rgb]{0.171875,0.2421875,0.3125}n_{p}$ learnable coefficients, and $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{P}_{1},\ldots,\mathbf{P}_{n_{p}}\in\mathbb{R}^{D_{emb}}$ are vectors that represent the teleconnection patterns, which typically can be obtained by manual extraction or automatic learning from data. We project $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{c}$ into query space to create a teleconnection query $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{q}^{tel}=\mathbf{c}\mathbf{W}^{Q}\in\mathbb{R}^{d_{k}}$ and compute attention scores $\color[rgb]{0.171875,0.2421875,0.3125}b_{j}$ between this query and all latitude keys. These scores are added as a bias to the standard attention logits:
$b_{j}=\frac{1}{\sqrt{d_{k}}}(\mathbf{q}^{tel}\cdot\mathbf{K}_{j}),\enspace\tilde{A}_{ij}=\frac{1}{\sqrt{d_{k}}}(\mathbf{Q}_{i}\cdot\mathbf{K}_{j})+\lambda\cdot b_{j},$ (8)
where $\color[rgb]{0.171875,0.2421875,0.3125}\lambda$ is a hyperparameter controlling teleconnection influence. This mechanism biases attention toward active global teleconnection patterns. For instance, during an MJO event, the model can emphasize connections between the tropical Indian Ocean/western Pacific and remote affected regions zhang2005madden , empowering teleconnection-aware forecasting.
Cross-Scale Interaction
@@
3.4 S2S Forecasting
For each latitude $\color[rgb]{0.171875,0.2421875,0.3125}i$, we process its feature vector $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{z}^{final}_{i}=\mathbf{Z}_{final}(i,:)\in\mathbb{R}^{D_{emb}}$ as follows: $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{y}_{i}=\mathbf{W}^{(2)}\text{GELU}(\mathbf{W}^{(1)}\mathbf{z}^{final}_{i}+\mathbf{b}^{(1)})+\mathbf{b}^{(2)}$, where $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{W}^{(1)}\in\mathbb{R}^{D_{emb}\times 2D_{emb}}$ and $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{W}^{(2)}\in\mathbb{R}^{2D_{emb}\times(2CW)}$. The output $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{y}_{i}\in\mathbb{R}^{2CW}$ is reshaped into a tensor representing predictions for both horizons at all longitudes for this latitude. Stacking these across all latitudes yields the final forecast fields $\color[rgb]{0.171875,0.2421875,0.3125}\hat{\mathbf{Y}}^{(1)}_{t_{15}:t_{28}},\hat{\mathbf{Y}}^{(2)}_{t_{29}:t_{42}}\in\mathbb{R}^{C\times H\times W}$ for the two target periods. We train the model using the loss function:
$\mathcal{L}=\frac{1}{2\times C\times H\times W}\Bigl{(}\left\lVert\hat{\mathbf{Y}}^{(1)}_{t_{15}:t_{28}}-\mathbf{Y}^{(1)}_{t_{15}:t_{28}}\right\rVert_{2}^{2}+\left\lVert\hat{\mathbf{Y}}^{(2)}_{t_{29}:t_{42}}-\mathbf{Y}^{(2)}_{t_{29}:t_{42}}\right\rVert_{2}^{2}\Bigr{)}.$ (9)
@@ 3.4 S2S Forecasting
4 Experiments
Datasets
[…]
Evaluation Metric
Following existing works liucirt ; nathaniel2024chaosbench ; rasp2024weatherbench , we adopt three metrics: (1) Root Mean Squared Error (RMSE) with a latitude-based weighting scheme that accounts for Earth’s spherical geometry; (2) Anomaly Correlation Coefficient (ACC), which measures the pattern similarity between predicted and observed anomalies relative to climatology; and (3) Spectral Divergence (SpecDiv), a physics-based metric that quantifies the deviation between the power spectra of prediction and target in the spectral domain using Kullback-Leibler divergence principles kullback1951information . For RMSE and SpecDiv, lower values indicate better performance, while higher ACC values signify better prediction skill. The detailed mathematical formulations of these metrics are provided in Appendix B.1.
Baselines
To rigorously evaluate TelePiT, we benchmark against state-of-the-art forecasting models in two categories: (1) data-driven baselines, including FourCastNetV2 kurth2023fourcastnet , Pangu-Weather bi2023accurate , GraphCast lam2023learning , ClimaX nguyen2023climax , and CirT liucirt , following the standard S2S evaluation protocol of nathaniel2024chaosbench ; and (2) physics-based baselines of leading operational S2S systems: ECMWF molteni1996ecmwf , UKMO williams2015met , NCEP saha2014ncep , CMA wu2019beijing . We provide more details of these baselines in Appendix B.2.
Table 1: Performance comparison of TelePiT against data-driven models. Lower RMSE ($\downarrow$) indicates better accuracy, while higher ACC ($\uparrow$) indicates better correlation with ground truth.
@@ 4 Experiments
Table 2: Comparison of physical consistency using SpecDiv metric (lower is better $\downarrow$).
@@
Experimental setup

### v3

@@ 3.2.2. Physics-Informed Neural ODE
To incorporate physical prior knowledge of atmospheric dynamics (Kashinath et al., 2021), we evolve each frequency band using a physics-informed neural ODE. Drawing from fluid dynamics principles, we formulate an ODE in the latent space (Verma et al., 2024) that combines fundamental atmospheric processes with data-driven corrections to handle complex, unresolved dynamics.
@@
For each multi-scale component $\mathbf{X}^{(\ell)}=[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{H}]\in\mathbb{R}^{H\times D_{emb}}$, we define an ODE that evolves the latent state through time:
(11) $\frac{d\mathbf{x}_{i}}{dt}=\gamma\cdot\tanh(\mathbf{R}_{i})$
@@
(12) $\small\mathbf{R}_{i}=\underbrace{\boldsymbol{\nu}\odot(\mathbf{x}_{i+1}-2\mathbf{x}_{i}+\mathbf{x}_{i-1})}_{\text{diffusion}}+\underbrace{\boldsymbol{\mu}\odot\frac{\mathbf{x}_{i+1}-\mathbf{x}_{i-1}}{2}}_{\text{advection}}+\underbrace{\boldsymbol{f}}_{\text{forcing}}+\underbrace{\alpha\cdot\text{MLP}(\mathbf{x}_{i})}_{\text{neural correction}}$}
@@
  • Diffusion term: $\boldsymbol{\nu}\odot(\mathbf{x}_{i+1}-2\mathbf{x}_{i}+\mathbf{x}_{i-1})$ approximates the second-order diffusion operator along latitude, representing the meridional diffusion of atmospheric properties. The learnable parameter $\boldsymbol{\nu}\in\mathbb{R}^{D_{emb}}$ modulates diffusion strength for each feature dimension.
@@
  • Advection term: $\boldsymbol{\mu}\odot\frac{\mathbf{x}_{i+1}-\mathbf{x}_{i-1}}{2}$ approximates the first-order spatial derivative, representing meridional transport by mean flow. The parameter $\boldsymbol{\mu}\in\mathbb{R}^{D_{emb}}$ controls advection intensity.
@@
  • Forcing term: $\boldsymbol{f}\in\mathbb{R}^{D_{emb}}$ represents external forcing, analogous to radiative and diabatic processes in the climate system.
@@
  • Neural correction: $\alpha\cdot\text{MLP}(\mathbf{x}_{i})$ learns complex, nonlinear dynamics not captured by simplified physical terms. Parameter $\alpha$ balances the physical terms with data-driven corrections.
@@
For numerical stability and physical realism, we apply boundary conditions $\mathbf{x}_{0}=\mathbf{x}_{H+1}=0$ (zero-padding at the poles) and constrain the rate of change using a hyperbolic tangent function. $\gamma$ is a small scaling factor that moderates the rate of change to prevent numerical instabilities. We employ Euler with a fixed step size $\Delta t$:
(13) $\mathbf{x}_{i}(t+\Delta t)=\mathbf{x}_{i}(t)+\Delta t\cdot\frac{d\mathbf{x}_{i}}{dt}\Big{|}_{t}.$
We evolve each frequency band $\mathbf{X}^{(\ell)}$ independently from time $t=0$ to $t=T$ (where $T$ represents one forecast step in latent space), yielding evolved representations $\tilde{\mathbf{X}}^{(\ell)}$. This scale-dependent evolution reflects physical reality that atmospheric processes operate differently across scales, diffusion dominates at smaller scales while advection plays a more significant role at larger scales (Vallis, 2017).
3.3. Teleconnection-Aware Transformer
3.3.1. Teleconnection-Aware Self-Attention
@@
Teleconnections, statistically significant correlations between weather and climate events at distant locations, such as the North Atlantic Oscillation (NAO) and Madden-Julian Oscillation (MJO), are fundamental for accurate S2S forecasting (Wallace and Gutzler, 1981). We encode these global patterns through a novel teleconnection-aware self-attention mechanism.
@@
For each frequency band $\tilde{\mathbf{X}}^{(\ell)}\in\mathbb{R}^{H\times D_{emb}}$, we first compute standard queries, keys, and values: $\mathbf{Q}=\tilde{\mathbf{X}}^{(\ell)}\mathbf{W}^{Q}$, $\mathbf{K}=\tilde{\mathbf{X}}^{(\ell)}\mathbf{W}^{K}$, $\mathbf{V}=\tilde{\mathbf{X}}^{(\ell)}\mathbf{W}^{V}$, where $\mathbf{W}^{Q},\mathbf{W}^{K},\mathbf{W}^{V}\in\mathbb{R}^{D_{emb}\times d_{k}}$ are learnable projections and $d_{k}=D_{emb}/N_{h}$ for $N_{h}$ heads.
@@
(14) $\mathbf{c}=\sum_{j=1}^{n_{p}}\omega_{j}\mathbf{P}_{j},\quad\text{with}\quad\boldsymbol{\omega}=\text{softmax}(\mathbf{\bar{x}}\mathbf{W}^{p})\in\mathbb{R}^{n_{p}},$
where $\mathbf{\bar{x}}=\frac{1}{H}\sum_{i=1}^{H}\tilde{\mathbf{x}}_{i}$ is the global atmosphere state, $\mathbf{W}^{p}\in\mathbb{R}^{D_{emb}\times n_{p}}$ projects this state onto $n_{p}$ learnable coefficients, and $\mathbf{P}_{1},\ldots,\mathbf{P}_{n_{p}}\in\mathbb{R}^{D_{emb}}$ are vectors that represent the teleconnection patterns, which typically can be obtained by manual extraction or automatic learning from data. We project $\mathbf{c}$ into query space to create a teleconnection query $\mathbf{q}^{tel}=\mathbf{c}\mathbf{W}^{Q}\in\mathbb{R}^{d_{k}}$ and compute attention scores $b_{j}$ between this query and all latitude keys. These scores are added as a bias to the standard attention logits:
(15) $b_{j}=\frac{1}{\sqrt{d_{k}}}(\mathbf{q}^{tel}\cdot\mathbf{K}_{j}),\enspace\tilde{A}_{ij}=\frac{1}{\sqrt{d_{k}}}(\mathbf{Q}_{i}\cdot\mathbf{K}_{j})+\lambda\cdot b_{j},$
where $\lambda$ is a hyperparameter controlling teleconnection influence. This mechanism biases attention toward active global teleconnection patterns. For instance, during an MJO event, the model can emphasize connections between the tropical Indian Ocean/western Pacific and remote affected regions (Zhang, 2005), empowering teleconnection-aware forecasting.
3.3.2. Cross-Scale Interaction
@@
(16) $\hat{\mathbf{X}}^{(\ell+1)}=\text{LayerNorm}(\text{MLP}([\hat{\mathbf{X}}^{(\ell)},\hat{\mathbf{X}}^{(\ell+1)}])),$
where $\cdot$$\cdot$] denotes concatenation. Finally, we integrate information across all scales:
(17) $\mathbf{Z}_{final}=\frac{1}{L+1}\sum_{\ell=0}^{L}\hat{\mathbf{X}}^{(\ell)},$
yielding a representation that captures atmospheric dynamics across all scales.
3.4. S2S Forecasting
For each latitude $i$, we process its feature vector as follows:
(18) $\mathbf{z}^{final}_{i}=\mathbf{Z}_{final}(i,:)\in\mathbb{R}^{D_{emb}},$
(19) $\mathbf{y}_{i}=\mathbf{W}^{(2)}\text{GELU}(\mathbf{W}^{(1)}\mathbf{z}^{final}_{i}+\mathbf{b}^{(1)})+\mathbf{b}^{(2)},$
where $\mathbf{W}^{(1)}\in\mathbb{R}^{D_{emb}\times 2D_{emb}}$ and $\mathbf{W}^{(2)}\in\mathbb{R}^{2D_{emb}\times(2CW)}$. The output $\mathbf{y}_{i}\in\mathbb{R}^{2CW}$ is reshaped into a tensor representing predictions for both horizons at all longitudes for this latitude. Stacking these across all latitudes yields the final forecast fields $\hat{\mathbf{Y}}^{(1)}_{t_{15}:t_{28}},\hat{\mathbf{Y}}^{(2)}_{t_{29}:t_{42}}\in\mathbb{R}^{C\times H\times W}$ for the two target periods. We train the model using the loss function:
(20) $\mathcal{L}=\frac{1}{2CHW}\Bigl{(}\left\lVert\hat{\mathbf{Y}}^{(1)}_{t_{15}:t_{28}}-\mathbf{Y}^{(1)}_{t_{15}:t_{28}}\right\rVert_{2}^{2}+\left\lVert\hat{\mathbf{Y}}^{(2)}_{t_{29}:t_{42}}-\mathbf{Y}^{(2)}_{t_{29}:t_{42}}\right\rVert_{2}^{2}\Bigr{)}.$
@@ 3.4. S2S Forecasting
Table 1. Performance comparison of TelePiT against data-driven models. Lower RMSE ($\downarrow$) indicates better accuracy, while higher ACC ($\uparrow$) indicates better correlation with ground truth.
@@
Table 2. Comparison of physical consistency using SpecDiv metric (lower is better $\downarrow$).
@@
Table 3. Performance comparison of TelePiT against operational forecasting systems on RMSE ($\downarrow$), ACC ($\uparrow$), SepcDiv ($\downarrow$) metrics.
Metric Model Weeks 3-4 Weeks 5-6
z500 z850 t500 t850 z500 z850 t500 t850
RMSE ($\downarrow$)
UKMO 64.2965 43.1836 2.1440 2.5763 63.0383 45.6263 2.2743 2.3924
NCEP 68.4408 44.7869 2.2250 2.7422 71.3320 46.8340 2.3433 2.8331
ECMWF 63.6737 42.3491 2.0667 2.3910 68.1156 45.3303 2.1772 2.5110
CMA 69.8326 47.0562 2.5975 2.9181 71.3996 48.5830 2.6388 2.9334
TelePiT 48.6714 31.0823 1.6849 1.8739 47.7618 30.4056 1.6985 1.8949
ACC ($\uparrow$)
UKMO 0.9726 0.9294 0.9804 0.9768 0.9695 0.9225 0.9781 0.9767
NCEP 0.9694 0.9206 0.9790 0.9774 0.9669 0.9133 0.9766 0.9760
ECMWF 0.9724 0.9289 0.9815 0.9828 0.9677 0.9157 0.9794 0.9807
CMA 0.9679 0.9188 0.9746 0.9744 0.9660 0.9121 0.9739 0.9738
TelePiT 0.9848 0.9626 0.9882 0.9900 0.9852 0.9643 0.9881 0.9897
SpecDiv ($\downarrow$)
UKMO 0.1173 0.1142 0.2408 0.0989 0.1330 0.1267 0.2905 0.1182
NCEP 0.5567 0.4875 4.3344 0.6459 0.5598 0.4826 4.2610 0.6135
ECMWF 0.3976 0.0599 0.1229 0.2375 0.3093 0.0687 0.0575 0.1187
CMA 0.1612 0.1671 0.2725 0.0318 0.1589 0.1544 0.1277 0.0324
TelePiT 0.0180 0.0175 0.0150 0.0045 0.0161 0.0168 0.0184 0.0042
4. Experiments
Datasets. We evaluate TelePiT using the ERA5 reanalysis dataset (Hersbach et al., 2020), as processed by the ChaosBench benchmark (Nathaniel et al., 2024). This dataset provides global atmospheric variables at multiple pressure levels and at the surface, with a uniform spatial resolution of $1.5^{\circ}$, corresponding to a grid of $121\times 240$. Inspired by the previous study (Liu et al., 2025; Nguyen et al., 2024; Bi et al., 2023; Kurth et al., 2023; Lam et al., 2023), we include six pressure-level fields: geopotential ($z$), specific humidity ($q$), temperature ($t$), zonal wind ($u$), meridional wind ($v$), and vertical velocity ($w$), each provided at ten discrete pressure levels ($10$, $50$, $100$, $200$, $300$, $500$, $700$, $850$, $925$, and $1000hPa$). Additionally, three single-level surface variables are included: 2-meter temperature ($t2m$), 10-meter u-wind ($10u$) and 10-meter v-wind ($10v$), yielding a total of 63 variables. We use data from 1979–2016 for training, 2017 for validation, and 2018 for testing.
Evaluation Metric. Following existing works (Liu et al., 2025; Nathaniel et al., 2024; Rasp et al., 2024), we adopt three metrics: (1) Root Mean Squared Error (RMSE) with a latitude-based weighting scheme that accounts for Earth’s spherical geometry; (2) Anomaly Correlation Coefficient (ACC), which measures the pattern similarity between predicted and observed anomalies relative to climatology; and (3) Spectral Divergence (SpecDiv), a physics-based metric that quantifies the deviation between the power spectra of prediction and target in the spectral domain using Kullback-Leibler divergence principles (Kullback and Leibler, 1951). For RMSE and SpecDiv, lower values indicate better performance, while higher ACC values signify better prediction skill. The detailed mathematical formulations of these metrics are provided in Appendix B.1.
Baselines. To rigorously evaluate TelePiT, we benchmark against state-of-the-art forecasting models in two categories: (1) data-driven baselines, including FourCastNetV2 (Kurth et al., 2023), Pangu-Weather (Bi et al., 2023), GraphCast (Lam et al., 2023), ClimaX (Nguyen et al., 2023), and CirT (Liu et al., 2025), following the standard S2S evaluation protocol of (Nathaniel et al., 2024); and (2) physics-based baselines of leading operational S2S systems: ECMWF (Molteni et al., 1996), UKMO (Williams et al., 2015), NCEP (Saha et al., 2014), CMA (Wu et al., 2019). We provide more details of baselines in Appendix B.2.
Figure 2. RMSE comparison on variable $z$, $t$, $u$, and $v$ of different pressure levels.